# Floquet breathers in a modulated nonlinear lattice

### v3

@@ V.A Period doubling $\Omega=2\omega_{\text{b}}$
The subject of time-modulation is of recent interest in photonic topological insulators. For example, Floquet solitons, for which the intensity repeats after each driving period, were found experimentally in optical waveguide arrays with Kerr nonlinearity Mukherjee and Rechtsman (2020) and period-doubled Floquet solitons were observed in a photonic topological insulator formed by an array of waveguides with a honeycomb pattern.Mukherjee and Rechtsman (2023) For these solutions, the intensity period is doubled. The same happens in a helically modulated honeycomb lattice as well and a sinusoidally driven SSH lattice. These references show that the period-doubling phenomenon can appear in many different physical systems of interest and, in particular, photonic systems are ideal experimental devices to observe the connection between time modulation, topology and nonlinearity.
Period doubling also appears in our system. In Fig. 5-bottom we can see the evolution of the breather center and its neighboring sites, also the evolution of the energy, which is $T_{m}$-periodic and therefore also $T_{\text{b}}$-periodic. This is a consequence of $u_{n}(T_{\text{b}}/2)=-u_{n}(0)$, $p_{n}(T_{\text{b}}/2)=-p_{n}(0)=0$, the energy being a function of $u_{n}^{2}$ and $\cos(\Omega t)=1$ at values of $t$ that are integer multiples of $T_{b}/2$. There is no net gain or loss of energy every $T_{\text{b}}/2$ with a relative precision of $10^{-9}$ after 100 breather periods using a symplectic integrator.Sanz-Serna and Calvo (1994)
@@ V.B Cases $\omega_{\text{b}}=3\Omega$ to $\omega_{\text{b}}=10\Omega$
For $\omega_{\text{b}}=3\Omega$, the breather is slightly unstable with two real multipliers very close to +1. In spite of that, simulations can be done to simulate the breather evolution with relatively long time. The Floquet multipliers are represented in Fig. 7 and the Fourier spectrum in a system with noise and $N=160$ oscillators is also shown.
For $\omega_{\text{b}}=n_{T}\Omega$, width $n_{T}=1\dots 10$, we have constructed breathers, for $n_{T}=3,4,8$, breathers are unstable, while the rest are stable, therefore there is no apparent pattern. For example, for $n_{T}=3$, there are two real eigenvalues slightly separated from +1+1, but, in spite of that, the breather has a long life. In Sec. VII, there is also more information on the stability of this type of breathers and its evolution with the frequency.
@@ VI Breathers with $\Omega=\omega_{\text{b}}/n_{T}$
A peculiarity of this case is that the linear stability of the breather and the Floquet matrix has to be calculated for a time equal to the modulation period $T_{m}=n_{T}T_{\text{b}}$ as the coordinates, and velocities have to repeat, but also the forces which depend on time and only repeat after $T_{m}$. Therefore, the phonon band appear with Floquet exponents multiplied by $n_{T}$ and spread out as they are given by $\theta_{q}=\omega_{q}T_{m}=n_{T}\omega_{q}T_{\text{b}}$ for a phonon of frequency $\omega_{q}$.
@@
In this section we separate the study of the Sievers-Takeno mode Sievers and Takeno (1988) or single breather, from the Page mode Page (1990) or double breather. Values $n_{T}=1,2,\dots,32$ are considered and breathers obtained. For $n_{T}=32,\dots,40$ phonobreathers Marín and Aubry (1996) appear as explained in Sec. VI.B.
@@ VI.A Sievers-Takeno mode: single ILM with $\omega_{\text{b}}=n_{T}\Omega$
We construct the ILM in a small lattice with $N=16$ cantilevers, and check its linear stability by constructing the monodromy matrix and obtaining the Floquet multipliers and exponents.Aubry (2006) Then, we embed the initial positions within a larger lattice with $N=160$ particles with some small random noise velocities, let it evolve and perform the XTDFT to check the results. An example $\Omega=\omega_{\text{b}}/10$ with $\omega_{\text{b}}=0.98$ can be seen in Fig. 8. The top panel shows the profile at $t=0$ and the Floquet eigenvalues, showing its stability. The bottom panel illustrates the evolution of the variables for a period of the AC current.
@@
Figure 10: (Top) For $\Omega<\omega_{\text{b}}/33$, a phonobreather appears with a background wave to the ILM. The inset shows the Floquet eigevalues with the Krein signature coded blue(+), red(-), black (zero). (Bottom) Fourier spectra. See text. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
Figure 9-top-right shows the XTDFT of the same breather but located within a larger system with $N=160$ and thermalized with initial random velocities. The breather survives and the phonon band becomes apparent. Two replicas of the phonon band can be seen with frequencies $\omega=\pm\Omega+\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q)}$, and also two replicas of the breather band with frequencies $\omega_{\text{b}}\pm\Omega$. Each breather-line replica is located below the corresponding phonon band replica. Other replicas with weaker intensity are out of the frame.
@@
Figure 11: (Top) For $\Omega<\omega_{\text{b}}/10$, profile of the unstable in-phase Page mode. (Bottom) The stable anti-phase Page mode with the same parameters and modulating frequency. The insets show the Floquet eigevalues with the Krein signature coded blue(+), red(-), black (zero). Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
@@ VI.B Phonobreathers: single ILM with phonon background for $\omega_{\text{b}}=n_{T}\Omega$
Although obtaining an ILM for increasing $n_{T}$ is possible, as $\Omega$ becomes smaller, the upper secondary ILM band $\omega_{\text{b}}+\Omega$ collides with the upper frequency of the phonon band $\omega_{\text{max}}=\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(\pi))}=\sqrt{\omega_{0}^{2}+4\kappa}=\omega_{\text{b}}+\Omega$, or $\Omega=\omega_{\text{max}}-\omega_{\text{b}}$ and $\omega_{\text{b}}/\Omega\simeq 32.3$, for the values $\omega_{0}=1$ and $\kappa=0.0052$. Then for $n_{T}>30$ the upper phonons are excited and a extended wave with $q=\pi$ profile appears as the background of the ILM, with smaller but comparable amplitude, forming a phonobreather. A phonobreather is a periodic solution on a nonlinear system formed by a breather with a background of homogeneous amplitude. They were first described by Marín and Aubry.Marín and Aubry (1996); Aubry (1997); Marín and Aubry (1998) The background is not a phonon as it is not a solution of the linearized system and has a significant amplitude, it can also be described as a nonlinear extended wave.
@@ VI.C Page modes with $\omega_{\text{b}}=n_{T}\Omega$
A Page mode Page (1990); Flach and Gorbach (2008) is a double breather or ILM with two sites vibrating with the same amplitude, it can be bond-symmetric if both sites oscillate in-phase (iP), or bond-antisymmetric if they oscillate in anti-phase (aP). The iP mode is unstable and the aP is stable for an unmodulated Klein Gordon system with soft potential.Marín et al. (1998) Both two modes appear also in the modulated system with the same stability properties. Both mode profiles and Floquet multipliers are represented in Fig. 11 for $\omega_{\text{b}}=10\Omega$. The XTDFT of the in-phase mode shows the appearance of another breather band closer to the bottom of the phonon band, which correspond to the unstable mode. Similar phenomena to the single ILM occur, i.e., formation of replicas of the phonon band and ILM band displaced $\pm\Omega$ in the frequency space and the excitation of the background forming a phonobreather when the secondary ILM band interacts with another replica of the phonon band just below.
@@ VII Dependence of the breather stability with the frequency
In this section, we observe the changes of stability and instability when both the breather frequency $\omega_{\text{b}}$ and $\Omega$ change simultaneously as they are commensurate as deduced before in Sec. IV. The number of possible cases, i.e, depending of the fraction $\omega_{\text{b}}/\Omega$ is huge, and the stability changes for each case are difficult to analyze. A panorama of the many cases is shown in Fig. 12, where a large number of cases for $\omega_{\text{b}}=(m_{\text{b}}/m_{\text{m}})\Omega$, with $m_{\text{b}}=10$ and $m_{\text{m}}=[2,\dots,30]$ integers are presented indicating their stability or instability type. Note that those integers are not necessarily the same as in Lemma 1 and 2 in Sec. IV.A, but their ratio is the same.
@@
Each breather calculation is represented by a circle with a color coding its instability as indicated by the legend. The breather circles make apparent lines with constant slope $\omega_{\text{b}}/\Omega=m_{\text{b}}/m_{\text{n}}$. A particularly interesting case are equal-period breathers with $\omega_{\text{b}}=\Omega$, where it can be seen that they can be continued into the phonon band. Other case of interest are period-doubled breathers, $\omega_{\text{b}}=\Omega/2$, for which there is a resonance with the negative phonon band $\omega_{q}^{(-)}=-\omega_{q}=-\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q))}$, which is equivalent in a system with modulation frequency $\Omega$ to $\Omega-\omega_{q}$ and the two phonon bands collide. As there is no dissipation the phonons take energy from the driving, grow and become nonlinear with their frequencies separating from the unmodulated phonon band.
@@
When observing that a breather is unstable, that may be due to interaction with phonons and other causes. A frequent one is that there is another breather with different structure which is stable for the same parameters. We tested for the existence and stability of breathers in quadrature, that is, with a phase difference of $\pi/2$ with the modulating term, with interesting results. Often, but by no means always, there is an interchange of stability between the in-phase breathers and the in-quadrature breathers as can be seen in Fig. 12.
@@
In this section we concentrate into three cases of interest for ST breathers (i) equal-period breathers, with $\omega_{\text{b}}=\Omega$, period-doubled breathers $\omega_{\text{b}}=\Omega/2$, and period-halving breathers $\omega_{\text{b}}=2\Omega$. All of them have been found experimentally in photonic systems Mukherjee and Rechtsman (2023, 2020); Kang et al. (2025) and the comparison may be of interest.
@@
In order to present the stability evolution, a very useful method is the simultaneous representation of the moduli and arguments of the Floquet multipliers $|\lambda_{i}|$ and $\arg{\lambda_{i}}$. Usually, the the existence of a breather appears as a pair of isolated eigenvalues corresponding to internal modes of the breather, and the phonons as a group or band of eigenvalues, and, of course, this simple picture can change a lot. We think it is more practical to defer most of the explanation to the figure captions and give here some brief indications.
@@
For period-doubled breathers the opposite happens: the in-phase breather is stable becoming unstable when crossing the phonon band but becoming stable again until the system itself becomes unstable as the two phonon bands collide. The in-quadrature breather is unstable but can be continued until the collision of the phonon bands. This properties can be seen in Fig. 14.
For period-halved breathers, the in-quadrature breather is again stable and the in-phase ones are unstable. The plots in this cases are trivial and not represented.
@@ VIII Conclusions
The nonlinear dynamical equations are derived, and is linear limit discussed. The dispersion bands, and its changes due to space, time, and space-time modulation have been obtained analytically and numerically.
We have adapted the theory of exact breathers to time-modulated systems, concluding that the modulation and breather frequencies in the moving frame should be commensurate as a necessary condition for breather existence.
The peculiarities of the stability analysis for time-modulated systems have been also analyzed, with the construction of an extended symplectic system.
First, breathers with a frequency multiple of the modulating one are studied and their properties analyzed. For modulating frequencies smaller than the breather frequency, single breathers and double breathers are found, both symmetric and anti-symmetric. When the modulating frequency becomes small a phonobreather with extended background appears.
@@
Some theoretical results on moving breathers in modulated systems presented in this article will be numerically studied and checked in a forthcoming article. This study reveals that this relatively simple system provides a huge variety of Floquet breathers which may also appears in other space-time modulated materials also known as dynamic metamaterials.
@@ Acknowledgements
All the authors would like to acknowledge Serge Aubry for the inspiration of his work. JFRA, in particular, although never worked directly with him, he did so with close collaborators as Robert Mackay and José Luis Marín in Cambridge, in 1997, where discussions with them over common publications with Aubry became his introduction to research in this field.

### v4

@@ V.A Period doubling $\Omega=2\omega_{\text{b}}$
The subject of time-modulation is of recent interest in photonic topological insulators. For example, Floquet solitons, for which the intensity repeats after each driving period, were found experimentally in optical waveguide arrays with Kerr nonlinearity Mukherjee and Rechtsman (2020) and period-doubled Floquet solitons were observed in a photonic topological insulator formed by an array of waveguides with a honeycomb pattern.Mukherjee and Rechtsman (2023) For these solutions, the intensity period is doubled. The same happens in a helically modulated honeycomb lattice as well and a sinusoidally driven SSH lattice. These references show that the period-doubling phenomenon can appear in many different physical systems of interest and, in particular, photonic systems are ideal experimental devices to observe the connection between time modulation, topology, and nonlinearity.
Period doubling also appears in our system. In Fig. 5-bottom we can see the evolution of the breather center and its neighboring sites, also the evolution of the energy, which is $T_{m}$-periodic and, therefore, also $T_{\text{b}}$-periodic. This is a consequence of $u_{n}(T_{\text{b}}/2)=-u_{n}(0)$, $p_{n}(T_{\text{b}}/2)=-p_{n}(0)=0$, the energy being a function of $u_{n}^{2}$ and $\cos(\Omega t)=1$ at values of $t$ that are integer multiples of $T_{b}/2$. There is no net gain or loss of energy every $T_{\text{b}}/2$ with a relative precision of $10^{-9}$ after 100 breather periods using a symplectic integrator.Sanz-Serna and Calvo (1994)
@@ V.B Cases $\omega_{\text{b}}=3\Omega$ to $\omega_{\text{b}}=10\Omega$
For $\omega_{\text{b}}=3\Omega$, the breather is slightly unstable with two real multipliers very close to +1. In spite of that, simulations can be done to simulate the breather evolution for a relatively long time. The Floquet multipliers are represented in Fig. 7 and the Fourier spectrum in a system with noise and $N=160$ oscillators is also shown.
For $\omega_{\text{b}}=n_{T}\Omega$, width $n_{T}=1\dots 10$, we have constructed breathers, for $n_{T}=3,4,8$, breathers are unstable, while the rest are stable, therefore, there is no apparent pattern. For example, for $n_{T}=3$, there are two real eigenvalues slightly separated from +1+1, but, in spite of that, the breather has a long life. In Sec. VII, there is also more information on the stability of this type of breathers and its evolution with the frequency.
@@ VI Breathers with $\Omega=\omega_{\text{b}}/n_{T}$
A peculiarity of this case is that the linear stability of the breather and the Floquet matrix has to be calculated for a time equal to the modulation period $T_{m}=n_{T}T_{\text{b}}$ as the coordinates, and velocities have to repeat, but also the forces which depend on time and only repeat after $T_{m}$. Therefore, the phonon band appears with Floquet exponents multiplied by $n_{T}$ and spread out as they are given by $\theta_{q}=\omega_{q}T_{m}=n_{T}\omega_{q}T_{\text{b}}$ for a phonon of frequency $\omega_{q}$.
@@
In this section, we separate the study of the Sievers-Takeno mode Sievers and Takeno (1988) or single breather, from the Page mode Page (1990) or double breather. Values $n_{T}=1,2,\dots,32$ are considered and breathers obtained. For $n_{T}=32,\dots,40$, phonobreathers Marín and Aubry (1996) appear as explained in Sec. VI.B.
@@ VI.A Sievers-Takeno mode: single ILM with $\omega_{\text{b}}=n_{T}\Omega$
We construct the ILM in a small lattice with $N=16$ cantilevers, and check its linear stability by constructing the monodromy matrix and obtaining the Floquet multipliers and exponents.Aubry (2006) Then, we embed the initial positions within a larger lattice with $N=160$ particles with some small random velocities, let it evolve and perform the XTDFT to check the results. An example $\Omega=\omega_{\text{b}}/10$ with $\omega_{\text{b}}=0.98$ can be seen in Fig. 8. The top panel shows the profile at $t=0$ and the Floquet eigenvalues, showing its stability. The bottom panel illustrates the evolution of the variables for a period of the AC current.
@@
Figure 10: (Top) For $\Omega<\omega_{\text{b}}/33$, a phonobreather appears with a background wave to the ILM. The inset shows the Floquet eigenvalues with the Krein signature coded blue(+), red(-), black (zero). (Bottom) Fourier spectra. See text. Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
Figure 9-top-right shows the XTDFT of the same breather but located within a larger system with $N=160$ and thermalized with initial random velocities. The breather survives, and the phonon band becomes apparent. Two replicas of the phonon band can be seen with frequencies $\omega=\pm\Omega+\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q)}$, and, also, tow replicas of the breather band with frequencies $\omega_{\text{b}}\pm\Omega$. Each breather-line replica is located below the corresponding phonon band replica. Other replicas with weaker intensity are out of the frame.
@@
Figure 11: (Top) For $\Omega<\omega_{\text{b}}/10$, profile of the unstable in-phase Page mode. (Bottom) The stable anti-phase Page mode with the same parameters and modulating frequency. The insets show the Floquet eigenvalues with the Krein signature coded blue(+), red(-), black (zero). Parameters: $\kappa=0.0052$, $\delta_{1}=0.1067$, $\delta_{2}=0.6996$, $\omega_{\text{b}}=0.98$.
@@ VI.B Phonobreathers: single ILM with phonon background for $\omega_{\text{b}}=n_{T}\Omega$
Although obtaining an ILM for increasing $n_{T}$ is possible, as $\Omega$ becomes smaller, the upper secondary ILM band $\omega_{\text{b}}+\Omega$ collides with the upper frequency of the phonon band $\omega_{\text{max}}=\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(\pi))}=\sqrt{\omega_{0}^{2}+4\kappa}=\omega_{\text{b}}+\Omega$, or $\Omega=\omega_{\text{max}}-\omega_{\text{b}}$ and $\omega_{\text{b}}/\Omega\simeq 32.3$, for the values $\omega_{0}=1$ and $\kappa=0.0052$. Then for $n_{T}>30$ the upper phonons are excited and a extended wave with $q=\pi$ profile appears as the background of the ILM, with smaller but comparable amplitude, forming a phonobreather. A phonobreather is a periodic solution on a nonlinear system formed by a breather with a background of homogeneous amplitude. They were first described by Marín and Aubry.Marín and Aubry (1996); Aubry (1997); Marín and Aubry (1998) The background is not a phonon as it is not a solution of the linearized system and has a significant amplitude. It can also be described as a nonlinear extended wave.
@@ VI.C Page modes with $\omega_{\text{b}}=n_{T}\Omega$
A Page mode Page (1990); Flach and Gorbach (2008) is a double breather or ILM with two sites vibrating with the same amplitude. It can be bond-symmetric if both sites oscillate in-phase (iP), or bond-antisymmetric if they oscillate in anti-phase (aP). The iP mode is unstable and the aP is stable for an unmodulated Klein Gordon system with soft potential.Marín et al. (1998) Both two modes appear also in the modulated system with the same stability properties. Both mode profiles and Floquet multipliers are represented in Fig. 11 for $\omega_{\text{b}}=10\Omega$. The XTDFT of the in-phase mode shows the appearance of another breather band closer to the bottom of the phonon band, which corresponds to the unstable mode. Similar phenomena to the single ILM occur, i.e., formation of replicas of the phonon band and ILM band displaced $\pm\Omega$ in the frequency space and the excitation of the background forming a phonobreather when the secondary ILM band interacts with another replica of the phonon band just below.
@@ VII Dependence of the breather stability with the frequency
In this section, we observe the changes of stability and instability when both the breather frequency $\omega_{\text{b}}$ and $\Omega$ change simultaneously as they are commensurate as deduced before in Sec. IV. The number of possible cases, i.e., depending on the fraction $\omega_{\text{b}}/\Omega$ is huge, and the stability changes for each case are difficult to analyze. A panorama of the many cases is shown in Fig. 12, where a large number of cases for $\omega_{\text{b}}=(m_{\text{b}}/m_{\text{m}})\Omega$, with $m_{\text{b}}=10$ and $m_{\text{m}}=[2,\dots,30]$ integers are presented indicating their stability or instability type. Note that those integers are not necessarily the same as in Lemma 1 and 2 in Sec. IV.A, but their ratio is the same.
@@
Each breather calculation is represented by a circle with a color coding its instability as indicated by the legend. The breather circles make apparent lines with constant slope $\omega_{\text{b}}/\Omega=m_{\text{b}}/m_{\text{n}}$. A particularly interesting case are equal-period breathers with $\omega_{\text{b}}=\Omega$, where it can be seen that they can be continued into the phonon band. Other case of interest are period-doubled breathers, $\omega_{\text{b}}=\Omega/2$, for which there is a resonance with the negative phonon band $\omega_{q}^{(-)}=-\omega_{q}=-\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q))}$, which is equivalent in a system with modulation frequency $\Omega$ to $\Omega-\omega_{q}$ and the two phonon bands collide. As there is no dissipation, the phonons take energy from the driving, grow and become nonlinear with their frequencies separating from the unmodulated phonon band.
@@
When observing that a breather is unstable, that may be due to interaction with phonons and other causes. A frequent one is that there is another breather with different structure which is stable for the same parameters. We tested the existence and stability of breathers in quadrature, that is, with a phase difference of $\pi/2$ with the modulating term, with interesting results. Often, but by no means always, there is an interchange of stability between the in-phase breathers and the in-quadrature breathers as can be seen in Fig. 12.
@@
In this section, we concentrate into three cases of interest for ST breathers (i) equal-period breathers, with $\omega_{\text{b}}=\Omega$, period-doubled breathers $\omega_{\text{b}}=\Omega/2$, and period-halving breathers $\omega_{\text{b}}=2\Omega$. All of them have been found experimentally in photonic systems Mukherjee and Rechtsman (2023, 2020); Kang et al. (2025) and the comparison may be of interest.
@@
In order to present the stability evolution, a very useful method is the simultaneous representation of the moduli and arguments of the Floquet multipliers $|\lambda_{i}|$ and $\arg{\lambda_{i}}$. Usually, the existence of a breather appears as a pair of isolated eigenvalues corresponding to internal modes of the breather, and the phonons as a group or band of eigenvalues, and, of course, this simple picture can change a lot. We think it is more practical to defer most of the explanation to the figure captions and give here some brief indications.
@@
For period-doubled breathers the opposite happens: the in-phase breather is stable becoming unstable when crossing the phonon band but becoming stable again until the system itself becomes unstable as the two phonon bands collide. The in-quadrature breather is unstable but can be continued until the collision of the phonon bands. These properties can be seen in Fig. 14.
For period-halved breathers, the in-quadrature breather is again stable, and the in-phase ones are unstable. The plots in this cases are trivial and not represented.
@@ VIII Conclusions
The nonlinear dynamical equations are derived, and its linear limit is discussed. The dispersion bands, and their changes due to space, time, and space-time modulation have been obtained analytically and numerically.
We have adapted the theory of exact moving breathers to space-time modulated systems, concluding that the modulation and breather frequencies in the moving frame should be commensurate as a necessary condition for breather existence. For stationary breathers, the condition applies to laboratory frequencies.
The peculiarities of the stability analysis for time-modulated systems have have also been analyzed with the construction of an extended symplectic system.
First, breathers with a frequency multiple of the modulating one are studied and their properties analyzed. For modulating frequencies smaller than the breather frequency, single breathers and double breathers are found, both symmetric and anti-symmetric. When the modulating frequency becomes small, a phonobreather with extended background appears.
@@
Some theoretical results on moving breathers in modulated systems presented in this article will be numerically studied and checked in a forthcoming article. This study reveals that this relatively simple system provides a huge variety of Floquet breathers which may also appear in other space-time modulated materials also known as dynamic metamaterials.
@@ Acknowledgements
All the authors would like to acknowledge Serge Aubry for the inspiration of his work. JFRA, in particular, although he never worked directly with him, he did so with close collaborators as Robert Mackay and José Luis Marín in Cambridge, in 1997, where discussions with them over their common publications with Aubry became his introduction to research in this field.